OpenStudy (anonymous):

Double Integrals y'all! ∫3 to 1 ∫1 to 0 (2xy)/((x^2)+1) dx dy

OpenStudy (across):

\[\int_{1}^{3}\int_{0}^{1}\frac{2xy}{x^2+1}dxdy\]What's so tough about this one?

OpenStudy (anonymous):

just do it in turns...

OpenStudy (anonymous):

first x, later y

OpenStudy (anonymous):

I'm studying for finals and going through my entire textbook and after having 4 other finals prior to today, you could say my brain is fried. sorry for being a college student and having off days

OpenStudy (across):

We know how you feel. All of us have gone through that.

OpenStudy (anonymous):

well kind of a harsh response before then.

OpenStudy (amorfide):

not harsh just saying it is easy to do, if you are in the right state of mind or if you get a quick tip on how to start it off

OpenStudy (amorfide):

let's not argue, just answer his/her question people :D

OpenStudy (anonymous):

I'm not here to argue. I just want to make sure I am doing things right, this is supposed to be study help yeah?

OpenStudy (amorfide):

this will explain better than we can

OpenStudy (amorfide):

limit of 0-1 is for x so integrate sub in 1 for x integrate again sub in y values sub in 3 - sub in values of 1 enjoy :D

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\int\limits_{0}^{1}(2xy /x ^{2}+1)dxdy=\int\limits_{1}^{3}y[Ln(x ^{2}+1]_{0}^{1}dy=Ln2/2(y ^{2})|_{1}^{3}=(Ln(2)/2)(9-1)=4Ln(2)\]

OpenStudy (anonymous):

lol, didn't fit....... till the end

OpenStudy (anonymous):

4Ln(2)

OpenStudy (anonymous):

Thanks everyone. I appreciate not allowing me to completely brainfart today